# Enhancing predictive capabilities in fusion burning plasmas through surrogate-based optimization in core transport solvers

### 19 Dec 2023 v1

@@ 1 Introduction
The local neoclassical and gyrokinetic calculations are then embedded into global transport solvers that compute self-consistent particle (fueling), momentum (torque) and energy (heating) sources to be balanced against the neoclassical and gyrokinetic losses, iterating until steady-state conditions are found. Even with the many simplifications provided by this embedded approach, the stiff behavior of turbulence results in the requirement of running many transport iterations to achieve steady-state (flux-matching) conditions. While it is hard to determine a priori how many flux-tube $\delta f$ simulations are required to simulate a given plasma condition, it is often the case that hundreds or even thousands of evaluations per radial location are required. This number varies from one case to another, but generally it increases with the number of channels simulated (e.g. simultaneous prediction of temperatures and densities is more expensive than only temperatures prediction), the evolution of targets (e.g. self-consistent alpha heating introduces a nonlinearity) and the stiffness, discontinuities and non-monotonicity of the transport model. In burning plasmas where heating is dominated by the profiles themselves via alpha heating, where density peaking comes mostly from transport processes and not sources, and with large gyro-Bohm units, the convergence of transport solvers becomes exceptionally difficult. Beginning with TGYRO-GYRO Candy2009a and TRINITY-GS2 Barnes2010 , efforts in the community are ongoing to develop advanced transport solvers (including TANGO-GENE siena_global_2022 and TRINITY-GX mandell_gx_2022 ) with computationally efficient transport models that capture the physics of gyrokinetic turbulence brought to macroscopic steady-state conditions.
@@ 2 Background
Eq. 1 clearly shows how coupled the system of FSA equations is. Every transport flux component ($\mathbf{\Gamma}_{e}$, $\mathbf{q}_{e}$, $\mathbf{q}_{i}$, $\mathbf{\Pi}$, $\mathbf{\Gamma}_{Z}$), depends on the state of background turbulence, which is in turn determined by each of the kinetic profiles under consideration. Furthermore, target fluxes such as energy exchange ($\langle P_{ei}\rangle$), radiation ($\langle P_{rad}\rangle$) and alpha power ($\langle P_{\alpha,e}\rangle$, $\langle P_{\alpha,i}\rangle$) strongly depend on the profiles themselves. Here it is assumed that the momentum flux, $\langle\mathbf{\Pi}\cdot\nabla r\rangle$, includes the residual stress, and that there are not intrinsic sources of the impurity of interest (hence the null-flux condition).
The process to bring an initial set of kinetic profiles to steady-state conditions when external heating, fueling and torque sources are imposed, requires a number of local, $\delta f$ turbulence simulations to provide the transport fluxes at each iteration.
Both in time dependent (with a criterion that accounts for the stationarity of the predicted quantities) and time independent (with a criterion based on the closeness between transport and target fluxes) solvers, the information of previous simulations is often not used, including the local Jacobian at each iteration. The paradigm presented here proposes that a global surrogate model for each of the transport fluxes is constructed and refined during the convergence process. All the simulations, including the ones far from convergence, are utilized to train the surrogate model and build the dependencies of transport fluxes with respect to the free parameters of the problem.
@@ 3 Flux-matching as a surrogate-based optimization problem
Figure 1: Visualization of problem geometry and free parameters of a three-channel prediction. a) Illustration of nested flux surfaces for a DIII-D ITER Similar Shape plasma howard_2024_pop , with magnetic axis in dashed black line, 5 selected flux surfaces in blue (cross section in purple) and last closed flux surface in green. Note that the lower x-point region is smoothed out due to Fourier-moments decomposition typical of transport solvers such as TRANSP Breslau2018 . Subplots b), d) and f) depict the piece-wise linear function representations of the normalized logarithmic gradient of electron temperature, ion temperature and electron density with the value at 5 selected flux surfaces (values to be predicted by the transport solver). Subplots c), e) and g) depict the corresponding kinetic profiles of each channel, produced by radial integration of logarithmic gradient from an edge boundary condition (dark blue).
@@ 3.1.2 Outcome space transformation
We employ a Monte-Carlo based objective function to transform the posterior distributions of each local channel residual into the scalarized $\xi$ total residual. By leveraging BoTorch’s sample-average approximation approach and PyTorch’s auto-differentiation capabilities, we are able to back-propagate gradients through the acquisition function, objective, GP model, and transforms all the way back to the inputs, which allows us to effectively optimize the acquisition function to determine the next simulation parameters for the expensive model.
@@ 3.1.3 Simple relaxation as initialization method
Instead of random training (employed in the first PORTALS publications rodriguez-fernandez_nonlinear_2022 ; howard_2024_pop ), we utilize the derivative-free simple relaxation (SR) method also implemented in TGYRO Candy2009a . The SR technique consists of neglecting cross-channel, cross-radius interactions and assuming that the local transport matrix is diagonal, with positive diffusion coefficients built from the relative difference between target and transport fluxes at each iteration. Formally, the SR iteration scheme can be written as:
@@
Figure 2: Performance of PORTALS at achieving 10×10$\times$10 × and 100×100$\times$100 × residual reduction with TGLF in two example H-mode plasmas: DIII-D ITER Similar Shape (ISS) and ITER Baseline howard_2024_pop . Random and SR initialization methods are compared, starting from the same initial condition. (top) Residual vs evaluations, with vertical line delimiting initialization phase and purple horizontal lines indicating the 10×10$\times$10 × and 100×100$\times$100 × residual reductions. (bottom) Violin plots representing distribution of number of evaluations that were required to achieve residual goals. Both methods were initialized with a random seed, and distributions were obtained for 16 seeds. As expected, SR is not strongly affected by the seed, particularly at the beginning of the flux matching process.
@@ 3.1.4 Physics-informed GP and acquisition optimization
In the current implementation of PORTALS, we utilize a Radial Basis Function (RBF) covariance kernel for the Gaussian processes of turbulent and neoclassical fluxes. We found that the use of a linear mean for gradient quantities, $z_{j,\forall c}$, and constant mean for profile quantities, $\widehat{y}_{j,\forall c}$, works well. This choice is motivated by the physics understanding that logarithmic gradients most strongly affect the background turbulence and profile quantities provide smaller corrections. This will be particularly true as the plasma reaches steady-state, a situation where profile quantities will barely change and turbulence will be above critical-gradient, thus a higher-order polynomial makes sense for $z_{j,\forall c}$. Future work will further utilize physics information to build better surrogates, to include constraints such as $F^{m}_{j,c}(z_{j,c}=0)=0$ (zero flux at zero gradient) or $\frac{\partial F^{m}_{j,c}}{\partial z_{j,c}}>0$ (positive diagonal of transport matrix). The addition of such constraints could help build better surrogate models that can accurately capture the system’s behavior with lower number of evaluations.
@@
In this work, we leverage the posterior mean of the predicted distribution of the residual $\xi$ to decide the next points to evaluate, which can be readily implemented with multi-dimensional root-finding methods (described below). In practice, other acquisition functions such as Expected Improvement (EI) (mockus_bayesian_1975, ; jones_efficient_1998, ), Upper Confidence Bound (UCB), or information-based variants such as entropy search (wang_max-value_2018, ) are often used in the Bayesian optimization literature, as they balance exploration (sampling uncertain regions) and exploitation (sampling promising regions) more effectively. We will investigate the effect of different acquisition function choices in PORTALS, including any generalizations of the existing acquisition functions to the multi-residual structure of Eq. (3), as part of our future work.
@@
To optimize the acquisition function (i.e. solving for flux-matching in the mean of the surrogate posterior distribution), we employ a sequential combination of three numerical methods. Firstly, we take advantage of the decomposition of the total residual, $\xi$, into $\left(N_{c}\cdot N_{\rho}\right)$ individual local channel residuals. For this, we use the multi-dimensional root-finding methods available in SciPy Virtanen2020 . Specifically, the Levenberg-Marquardt (LM) method levenberg_method_1944 ; marquardt_algorithm_1963 has shown superior performance for this particular problem. The exact Jacobian of the surrogate models $F_{j,c}^{m}$ is provided to the LM method, as facilitated by automatic differentiation in PyTorch Paszke2019 , encompassing all transformations outlined in Sections 3.1.1 and 3.1.2 as well as the residual reconstruction in Eqs. 3. Secondly, in cases where the multi-variate LM method fails to achieve a satisfactory level of residual reduction, we resort to leveraging multi-start, scalarized-residual optimization in BoTorch. A heuristic generation of initialization points is used from which to start the L-BFGS-B local optimization algorithm Zhu1997 . Lastly, we utilize a genetic algorithm from the DEAP package Fortin2012 , to check for additional, possibly better optima that might have been missed by the local search algorithms of the previous steps.
@@ 3.1.5 Other physics considerations
As written in Eq. 1, the impurity density ($n_{Z}$) evolution is assumed to be source free, and therefore the workflow aims at solving for the null-flux condition, $\langle\mathbf{\Gamma}_{Z}\cdot\nabla r\rangle=0$. Impurity density peaking in such situations happens by means of a convective pinch that arises from neoclassical and turbulent transport and that results in a detrimental inward flux of impurities to the plasma core. Formulating the impurity density flux with a diffusion-convection ansatz, summing over charge states and ignoring the effect of charge on core transport [^4] We note that recent work howard_2024_pop suggests that this approximation to study impurity transport might need to be revisited and future work with PORTALS could include a charge-state ionization recombination model.:
@@
The framework of FSA transport equations described in Eq. 1 also allows the generalization of the energy exchange power, $\langle P_{ei}\rangle$, to include contributions from both classical and turbulent exchange. While the classical exchange can be calculated from standard net-exchange collisions formulas (see original PORTALS rodriguez-fernandez_nonlinear_2022 ), the turbulent exchange or anomalous heating is provided by the turbulence $\delta f$ simulations Candy2013 . This means that the turbulent exchange becomes an additional quantity to include during the residual calculation and requires its own surrogate model. At flux-matching, however, the turbulent exchange enters as a target flux density in Eqs. 3 and needs to be volume-integrated. This is the only component in the target calculations that cannot be extended to the finer radial grid described previously, and hence its effect on the macroscopic profiles will be subject to volume integration errors.
@@ 3.2 Benchmarking examples with the TGLF model
Figure 4: Comparison of residual evolution in PORTALS with random initial training, a standard Newton method (NM) and simple relaxation (SR) schemes. For this exercise, we employed the quasilinear Trapped Gyro-Landau Fluid (TGLF) SAT0 model Staebler2007 . Three example plasmas were used: (a) DIII-D ISS howard_2024_pop , (b) ITER Baseline howard_2024_pop and (c) JET Ohmic L-mode prf_2023_eps . PORTALS was initialized with different random seeds, which affect the initial training (5 profiles, vertical dashed line) and the subsequent training of Gaussian processes and acquisition optimization techniques. NM corresponds to TGYRO method-$1$ and each of the three numerical parameters (relaxation parameter $\eta$, maximum step size $b_{max}$ and finite differences step $\Delta_{x}$) were varied within reasonable ranges. SR corresponds to TGYRO method-$6$ with variations of the relaxation parameter $\eta_{SR}$. We note that, even though TGYRO works with gyro-Bohm normalized residuals, here we plot them in real units, $MW/m^{2}$, for a direct comparison of performance.
@@
Figure 5: Characteristic PORTALS phases during the prediction of three burning plasmas ([left] SPARC Primary Reference Discharge (PRD) Rodriguez-Fernandez2020a , [center] ARC holland_2023_aps and [right] ITER howard_2024_pop ) with nonlinear CGYRO. Phase I is the initial surrogate training phase, Phase II represents the first PORTALS iterations (often with oscillatory behavior as it learns key parametric dependencies) and Phase III is the final convergence phase. The transition from II to III is smooth and not characterized by a change in the PORTALS iteration scheme but it represents a transition to a situation where the surrogates are capturing the turbulence dynamics with enough accuracy to drive the system towards steady state.
Fig. 5 shows characteristic phases for the decrease of multi-channel residual of PORTALS-CGYRO simulations, for three examples. Phase I consists of the initial SR training, which often decreases residual if channels are not coupled or strongly dominated by one type of turbulence mode, although the simplicity of this scheme (as addressed in Sec. 3.1.3) can also lead to oscillations or higher residual (as in the middle plot in Fig. 5). However, the goal of this phase is not to observe a steady decrease of the residual, but of producing proper training points for subsequent phases of the PORTALS workflow. Phase II is when surrogates are fitted and acquisition function optimization informs next points to evaluate. During Phase II, oscillations are usually observed for 2-5 evaluations, a consequence of the surrogates exploring cases near the logarithmic gradients bounds, or combinations that were not explored during Phase I. In Phase III, the residual drops dramatically and flux-matching occurs. Sometimes, following Phase III, we observe oscillations due to critical gradient and stiff behaviors of turbulence. Profiles and gradients are converged, but fluxes can jump up and down the targets, as was discussed in Ref. rodriguez-fernandez_nonlinear_2022 and will be described in detail in Sec. 5.

### 09 Apr 2024 v2

@@ 1 Introduction
The local neoclassical and gyrokinetic calculations are then embedded into global transport solvers that compute self-consistent particle (fueling), momentum (torque) and energy (heating) sources to be balanced against the neoclassical and gyrokinetic losses, iterating until steady-state conditions are found. Even with the many simplifications provided by this embedded approach, the stiff behavior of turbulence results in the requirement of running many transport iterations to achieve steady-state (flux-matching) conditions. While it is hard to determine a priori how many flux-tube $\delta f$ simulations are required to simulate a given plasma condition, it is often the case that hundreds or even thousands of evaluations per radial location are required. This number varies from one case to another, but generally it increases with the number of channels simulated (e.g. simultaneous prediction of temperatures and densities is more expensive than only temperatures prediction), the evolution of targets (e.g. self-consistent alpha heating introduces a nonlinearity) and the stiffness, discontinuities and non-monotonicity of the transport model. In burning plasmas where heating is dominated by the profiles themselves via alpha heating, where density peaking comes mostly from transport processes and not sources, and with large gyro-Bohm units, the convergence of transport solvers becomes exceptionally difficult. Beginning with TGYRO-GYRO Candy2009a and TRINITY-GS2 Barnes2010 whose pioneering studies set the foundation for work in this area, efforts in the community are ongoing to develop advanced transport solvers (including TANGO-GENE siena_global_2022 and TRINITY-GX mandell_gx_2022 ) with computationally efficient transport models that capture the physics of gyrokinetic turbulence brought to macroscopic steady-state conditions. The work presented here is complementary to these efforts, as it focuses on the acceleration of the convergence of transport solvers, rather than the development of the transport models themselves that are efficient and applicable to a wide range of plasmas (e.g. for global gyrokinetics and 3D geometry). The tokamak transport modeling community has also recently begun to employ machine learning-accelerated and data-driven models for transport predictions Meneghini2016 ; VanDePlassche2020 ; Ho2021 ; zanisi_efficient_2024 ; hornsby_gaussian_2024 . These efforts are primarily aimed at developing high-dimensional surrogate models to simulate transport fluxes and integrating them into standard transport solvers.
@@ 2 Background
Eq. 1 clearly shows how coupled the system of FSA equations is. Every transport flux component ($\mathbf{\Gamma}_{e}$, $\mathbf{q}_{e}$, $\mathbf{q}_{i}$, $\mathbf{\Pi}$, $\mathbf{\Gamma}_{Z}$), depends on the state of background turbulence, which is in turn determined by each of the kinetic profiles under consideration. Furthermore, target fluxes such as energy exchange ($\langle P_{ei}\rangle$), radiation ($\langle P_{rad}\rangle$) and alpha power ($\langle P_{\alpha,e}\rangle$, $\langle P_{\alpha,i}\rangle$) strongly depend on the profiles themselves. Here it is assumed that the momentum flux, $\langle\mathbf{\Pi}\cdot\nabla r\rangle$, includes the residual stress, and that there are not intrinsic sources of the impurity of interest (hence the null-flux condition). In the reminder of this paper, auxiliary heating power delivered to each species is assumed to be fixed throughout the simulations and therefore must be obtained by external means, such as interpretive simulations with TRANSP or ASTRA, and assumed constant with the variations of the kinetic profiles from initial to steady-state conditions. This is a current limitation of the framework, but it is expected to be addressed in future work. The rest of target flux components are calculated self-consistently with the kinetic profiles. Details on these calculations are available in Ref. rodriguez-fernandez_nonlinear_2022 .
The process to bring an initial set of kinetic profiles to steady-state conditions when external heating, fueling and torque sources are imposed, requires a number of local, $\delta f$ turbulence simulations to provide the transport fluxes at each iteration. Both in time dependent (with a criterion that accounts for the stationarity of the predicted quantities) and time independent (with a criterion based on the closeness between transport and target fluxes) solvers, the information of previous simulations is often not used, including the local Jacobian at each iteration. The paradigm presented here proposes that a global surrogate model for each of the transport fluxes is constructed and refined during the convergence process. All the simulations, including the ones far from convergence, are utilized to train the surrogate model and build the dependencies of transport fluxes with respect to the free parameters of the problem.
@@ 3 Flux-matching as a surrogate-based optimization problem
Figure 1: Visualization of problem geometry and free parameters of a three-channel prediction. a) Illustration of nested flux surfaces for a DIII-D ITER Similar Shape plasma howard_simultaneous_2024 , with magnetic axis in dashed black line, 5 selected flux surfaces in blue (cross section in purple) and last closed flux surface in green. Note that the lower x-point region is smoothed out due to Fourier-moments decomposition typical of transport solvers such as TRANSP Breslau2018 . Subplots b), d) and f) depict the piece-wise linear function representations of the normalized logarithmic gradient of electron temperature, ion temperature and electron density with the value at 5 selected flux surfaces (values to be predicted by the transport solver). Subplots c), e) and g) depict the corresponding kinetic profiles of each channel, produced by radial integration of logarithmic gradient from an edge boundary condition (dark blue).
@@
For clarification, we must note that the results obtained with PORTALS are not results of surrogate modeling of turbulence and transport, but of direct turbulence modeling (either with quasilinear gyrofluid TGLF or nonlinear gyrokinetic CGYRO for the cases shown in this paper). Surrogates are simply used to accelerate multi-channel convergence, but the steady-state predictions of profiles are a full-physics result and must be interpreted as such.
@@ 3.1.1 Input space transformation
We note that in the current implementation, separate surrogate models for turbulent and neoclassical fluxes are used, but the framework is flexible enough to include a single model for both, if the user is interested in a more general approach. In cases where the neoclassical fluxes are of high fidelity (e.g. with the NEO code belli_eulerian_2009 ), even if much cheaper than nonlinear gyrokinetic calculations, it is important to build surrogate models to their flux response. This is because their evaluation (unless analytical) will be much more expensive than a GP evaluation, and automatic differentiation would be, generally, not available; making the acquisition function optimization (the flux-matching process per se) significantly more expensive.
@@ 3.1.2 Outcome space transformation
Note that since the total residual $\xi$ in (3c) is a nonlinear function of the individual residuals $R_{j,c}$, its posterior distribution under the surrogate model is non-Gaussian. We therefore employ a Monte-Carlo based objective function to transform the posterior distributions of each local channel residual into the scalarized $\xi$ total residual. By leveraging BoTorch’s sample-average approximation approach and PyTorch’s auto-differentiation capabilities, we are able to back-propagate gradients through the acquisition function, objective, GP model, and transforms all the way back to the inputs, which allows us to effectively optimize the acquisition function to determine the next simulation parameters for the expensive model.
@@ 3.1.3 Simple relaxation as initialization method
Instead of random training (employed in the first PORTALS publications rodriguez-fernandez_nonlinear_2022 ; howard_simultaneous_2024 ), we utilize the derivative-free simple relaxation (SR) method also implemented in TGYRO Candy2009a . The SR technique consists of neglecting cross-channel, cross-radius interactions and assuming that the local transport matrix is diagonal, with positive diffusion coefficients built from the relative difference between target and transport fluxes at each iteration. Formally, the SR iteration scheme can be written as:
@@
Figure 2: Performance of PORTALS at achieving 10×10$\times$10 × and 100×100$\times$100 × residual reduction with TGLF in two example H-mode plasmas: DIII-D ITER Similar Shape (ISS) and ITER Baseline howard_simultaneous_2024 . Random and SR initialization methods are compared, starting from the same initial condition. (top) Residual vs evaluations, with vertical line delimiting initialization phase and purple horizontal lines indicating the 10×10$\times$10 × and 100×100$\times$100 × residual reductions. (bottom) Violin plots representing distribution of number of evaluations that were required to achieve residual goals. Both methods were initialized with a random seed, and distributions were obtained for 16 seeds. As expected, SR is not strongly affected by the seed, particularly at the beginning of the flux matching process.
@@ 3.1.4 Physics-informed GP and acquisition optimization
In the current implementation of PORTALS, we utilize a Radial Basis Function (RBF) covariance kernel for the Gaussian processes of turbulent and neoclassical fluxes. We found that the use of a linear mean for gradient quantities, $z_{j,\forall c}$, and constant mean for profile quantities, $\widehat{y}_{j,\forall c}$, works well for a few test cases performed so far. However, a proper characterization of the benefit of this approach v.s. the same polynomial degree for all input parameters has not been performed yet in a general manner and is subject of future work. This choice is motivated by the physics understanding that logarithmic gradients most strongly affect the background turbulence and profile quantities provide smaller corrections. This is expected to be particularly true as the plasma reaches steady-state, a situation where profile quantities will barely change and turbulence will be above critical-gradient, thus a higher-order polynomial makes sense for $z_{j,\forall c}$. Future work will further utilize physics information to build better surrogates, to include constraints such as $F^{m}_{j,c}(z_{j,c}=0)=0$ (zero flux at zero gradient) or $\frac{\partial F^{m}_{j,c}}{\partial z_{j,c}}>0$ (positive diagonal of transport matrix). The addition of such constraints could help build better surrogate models that can accurately capture the system’s behavior with lower number of evaluations.
@@
In this work, we leverage the posterior mean of the predicted distribution of the residual $\xi$ to decide the next points to evaluate, which can be readily implemented with multi-dimensional root-finding methods (described below). In practice, other acquisition functions such as Expected Improvement (EI) (mockus_bayesian_1975, ; jones_efficient_1998, ), Upper Confidence Bound (UCB), or information-based variants such as entropy search (wang_max-value_2018, ) are often used in the Bayesian optimization literature, as they balance exploration (sampling uncertain regions) and exploitation (sampling promising regions) more effectively. We will investigate the effect of different acquisition function choices in PORTALS, including any generalizations of the existing acquisition functions to the multi-residual structure of Eq. (3), as part of our future work. However, as noted in Section 3.1.2, as the total residual $\xi$ is a non-linear function of the individually modeled residuals, the full posterior distribution of these residuals is used in computing our target objective and thus the probabilistic nature of our surrogate models is crucial to our approach.
@@
To optimize the acquisition function (i.e. solving for flux-matching in the mean of the surrogate posterior distribution), we employ a sequential combination of three numerical methods. Firstly, we take advantage of the decomposition of the total residual, $\xi$, into $\left(N_{c}\cdot N_{\rho}\right)$ individual local channel residuals. For this, we use the multi-dimensional root-finding methods available in SciPy Virtanen2020 . Specifically, the Levenberg-Marquardt (LM) method levenberg_method_1944 ; marquardt_algorithm_1963 has shown superior performance for this particular problem. The exact Jacobian of the surrogate models $F_{j,c}^{m}$ is provided to the LM method, as facilitated by automatic differentiation in PyTorch Paszke2019 , encompassing all transformations outlined in Sections 3.1.1 and 3.1.2 as well as the residual reconstruction in Eqs. 3. We must note that this first optimization method operates directly using the mean of the prediction of each surrogate flux evaluation, and therefore does not take advantage of the Monte-Carlo based objective transformation that would properly capture non-Gaussian properties of the nonlinear Equation (3c). This would be required when uncertainty estimations from the Gaussian processes are taken into consideration for more advanced acquisition functions. This method was still kept in PORTALS as part of the historical familiarity of steady-state, flux-matching frameworks as multi-variable root-finding problems. Future work will be dedicated to understanding to what degree the decomposition of the total residual into each channel, radial flux components can be leveraged for acquisition optimization when moving away from the simplistic posterior-mean acquisition.
Secondly, in cases where the multi-variate LM method fails to achieve a satisfactory level of residual reduction, we resort to leveraging multi-start, scalarized-residual optimization in BoTorch. A heuristic generation of initialization points is used from which to start the L-BFGS-B local optimization algorithm Zhu1997 . Lastly, we utilize a genetic algorithm from the DEAP package Fortin2012 , to check for additional, possibly better optima that might have been missed by the local search algorithms of the previous steps.
@@ 3.1.5 Other physics considerations
As written in Eq. 1, the impurity density ($n_{Z}$) evolution is assumed to be source free, and therefore the workflow aims at solving for the null-flux condition, $\langle\mathbf{\Gamma}_{Z}\cdot\nabla r\rangle=0$. Impurity density peaking in such situations happens by means of a convective pinch that arises from neoclassical and turbulent transport and that results in a detrimental inward flux of impurities to the plasma core. Formulating the impurity density flux with a diffusion-convection ansatz, summing over charge states and ignoring the effect of charge on core transport [^4] We note that recent work howard_simultaneous_2024 suggests that this approximation to study impurity transport might need to be revisited and future work with PORTALS could include a charge-state ionization recombination model.:
@@
The framework of FSA transport equations described in Eq. 1 also allows the generalization of the energy exchange power, $\langle P_{ei}\rangle$, to include contributions from both classical and turbulent exchange. While the classical exchange can be calculated from standard net-exchange collisions formulas (see original PORTALS rodriguez-fernandez_nonlinear_2022 ), the turbulent exchange or anomalous heating is provided by the turbulence $\delta f$ simulations Candy2013 . This means that the turbulent exchange becomes an additional quantity to include during the residual calculation and requires its own surrogate model. At flux-matching, however, the turbulent exchange enters as a target flux density in Eqs. 3 and needs to be volume-integrated. This is the only component in the target calculations that cannot be extended to the finer radial grid described previously, and hence its effect on the macroscopic profiles will be subject to volume integration errors. We must clarify that the turbulent exchange or any other target flux that requires surrogate modeling (when non-analytic) is incorporated into the surrogate modeling framework via the target flux density, $f^{\text{target}}_{j,c}$, and not via the integrated flux. The volume integration is performed as part of the transformation from flux (surrogate) quantities to residual quantities during optimization.
@@ 3.2 Benchmarking examples with the TGLF model
Figure 4: Comparison of residual evolution in PORTALS with random initial training, a standard Newton method (NM) and simple relaxation (SR) schemes. For this exercise, we employed the quasilinear Trapped Gyro-Landau Fluid (TGLF) SAT0 model Staebler2007 . Three example plasmas were used: (a) DIII-D ISS howard_simultaneous_2024 , (b) ITER Baseline howard_simultaneous_2024 and (c) JET Ohmic L-mode prf_2023_eps . PORTALS was initialized with different random seeds, which affect the initial training (5 profiles, vertical dashed line) and the subsequent training of Gaussian processes and acquisition optimization techniques. NM corresponds to TGYRO method-$1$ and each of the three numerical parameters (relaxation parameter $\eta$, maximum step size $b_{max}$ and finite differences step $\Delta_{x}$) were varied within reasonable ranges. SR corresponds to TGYRO method-$6$ with variations of the relaxation parameter $\eta_{SR}$. We note that, even though TGYRO works with gyro-Bohm normalized residuals, here we plot them in real units, $MW/m^{2}$, for a direct comparison of performance.
@@
Figure 5: Characteristic PORTALS phases during the prediction of three burning plasmas ([left] SPARC Primary Reference Discharge (PRD) Rodriguez-Fernandez2020a , [center] ARC holland_2023_aps and [right] ITER howard_2024_iter ) with nonlinear CGYRO. Phase I is the initial surrogate training phase, Phase II represents the first PORTALS iterations (often with oscillatory behavior as it learns key parametric dependencies) and Phase III is the final convergence phase. The transition from II to III is smooth and not characterized by a change in the PORTALS iteration scheme but it represents a transition to a situation where the surrogates are capturing the turbulence dynamics with enough accuracy to drive the system towards steady state. Bottom subplots show the evolution of the residual per radial location separately, shaded from red (core) to blue (edge).
Fig. 5 shows characteristic phases for the decrease of multi-channel residual of PORTALS-CGYRO simulations, for three examples. Phase I consists of the initial SR training, which often decreases residual if channels are not coupled or strongly dominated by one type of turbulence mode, although the simplicity of this scheme (as addressed in Sec. 3.1.3) can also lead to oscillations or higher residual (as in the middle plot in Fig. 5). However, the goal of this phase is not to observe a steady decrease of the residual, but of producing proper training points for subsequent phases of the PORTALS workflow. Phase II is when surrogates are fitted and acquisition function optimization informs next points to evaluate. During Phase II, oscillations are usually observed for 2-5 evaluations, a consequence of the surrogates exploring cases near the logarithmic gradients bounds, or combinations that were not explored during Phase I. In Phase III, the residual drops dramatically and flux-matching occurs. Sometimes, following Phase III, we observe oscillations due to critical gradient and stiff behaviors of turbulence. Profiles and gradients are converged, but fluxes can jump up and down the targets, as was discussed in Ref. rodriguez-fernandez_nonlinear_2022 and will be described in detail in Sec. 5. Generally, as depicted in the bottom subplots in Fig. 5, the residual evolution is not uniform across radial locations, but no specific trends are found. In some situations, the core reaches low residuals faster than the edge, and vice versa.
Due to the need for expert knowledge to interpret if sufficient convergence has been achieved, together with the need for a careful nonlinear gyrokinetic setup (to avoid wasting computational resources), the current implementation of PORTALS requires significant “human in the loop” component. Future work will focus on the development of a more automated approach that can make high-fidelity profile predictiosn available to a wider user base.